### v3

# Mergers Fall Short: Non-merger Channels Required for Galactic Heavy Element Production

Abstract
Since the discovery of the binary neutron star merger GW170817 and its associated kilonova, neutron star mergers have been established as a key production channel for -process elements in the Universe. However, various evidences, including the observations of -process abundances as inferred from stellar spectra of Milky Way disk stars, suggest that additional channels are needed to fully account for the -process element enrichment in the Milky Way. Neutron star–black hole mergers and fast-merging binary neutron star systems are among the leading alternative candidates. In this paper, we combine gravitational-wave observations from LIGO–Virgo–KAGRA with data from short gamma-ray bursts, Galactic pulsars, and Galactic [Eu/Fe] versus [Fe/H] abundance observations to assess the contribution of these mergers to -process enrichment in the Galactic disk. Our analysis employs a unified, likelihood-based inference framework that consistently propagates uncertainties in merger rates, delay-time distributions, mass- and spin-dependent ejecta yields, and stellar abundance measurements. We find that neither neutron star–black hole mergers nor fast-merging binary neutron star populations can serve as the dominant additional channel without generating strong tension with existing observations and theoretical expectations. These results constrain the viable sources of Galactic -process enrichment and underscore the necessity of non-merger production channels.
I Introduction
The origins of the heaviest elements in the Universe via rapid neutron capture (-process) nucleosynthesis remains a key question in astrophysics. While it is known that the -process happens in neutron-rich environments Burbidge et al. (1957); Cameron (1957), their dominant astrophysical sites are still debated Cowan et al. (2021); Siegel (2022). Compact binary mergers, particularly binary neutron star (BNS) mergers, have been recognized as promising sources of -process elements Lattimer and Schramm (1974); Eichler et al. (1989); Freiburghaus et al. (1999). The detection of GW170817 and its associated kilonova provided direct evidence for -process nucleosynthesis in BNS mergers Abbott and Collaboration) (2017); Abbott et al. (2017); Kasen et al. (2017). However, the relative role of BNS mergers compared to other potential channels, such as neutron star-black hole (NSBH) mergers Rosswog (2005); Kyutoku et al. (2015), magnetorotational supernovae Winteler (2012); Nishimura et al. (2015), collapsars Siegel et al. (2019a), and giant flares from magnetars Cehula et al. (2024); Patel et al. (2025) is an open problem. For example, recent chemical evolution studies have raised concerns about whether the event rate evolution of BNS mergers are consistent with the early enrichment seen in metal-poor stars in the Milky Way Côté (2019); Hotokezaka et al. (2018); van de Voort et al. (2015); Cescutti et al. (2015); Kobayashi et al. (2020); van de Voort et al. (2022); Kobayashi et al. (2023); Kolborg et al. (2023). Similar concerns and conclusions are derived from chemical evolution studies of (ultra-faint) dwarf galaxies and globular clusters Bonetti et al. (2019); Zevin et al. (2019); Skúladóttir et al. (2019); Naidu et al. (2022); Simon et al. (2023); Kirby et al. (2023).
A key observational probe of -process enrichment in the Milky Way is the relationship between and (where ; denotes number abundances and denotes solar), which traces the evolution of europium (Eu, a representative -process element) relative to iron (Fe) as the Galaxy chemically evolves. Observations of metal-poor stars, particularly in the Galactic halo and thick disk, reveal a significant presence of Eu even at low Sneden et al. (2008); Frebel (2010), indicating that -process production was already occurring in the early history of the Galaxy. In order to match the early enrichment, at least some events must occur promptly after star formation—an argument supported by stars like J15213538, an extremely -process-enhanced star with and Cain et al. (2020). The relative abundance of Eu decreases at higher metallicities, implying an increase in late-time Fe production and/or decrease in Eu enrichment. These trends—particularly the negative slope in the [Eu/Fe] versus [Fe/H] relation observed in Milky Way disk stars—place stringent constraints on the rates and yields of -process sources throughout cosmic history. Here, by “yield” we mean the per-event -process yield, i.e. the total mass of -process material that a single event ejects into the Galactic interstellar medium (often referred to as the nucleosynthesis or chemical yield). Using – evolution as a reference, one can test whether the proposed astrophysical sites can reproduce the chemical evolution history of the Milky Way.
In our recent study Chen et al. (2025), we investigated whether BNS mergers alone can account for the observed -process enrichment of disk stars in the Milky Way ([Fe/H]). Standard BNS (“delayed BNS”) formation scenarios predict a delay relative to the star formation rate (SFR) Peters (1964); Beniamini and Piran (2019). Using a one-zone chemical evolution model Siegel et al. (2019b) informed by cosmic SFR models Madau and Fragos (2017), short GRB (sGRB)-inferred delay-time distributions Zevin et al. (2022), GW-inferred BNS rates and mass distributions Abbott et al. (2023a, b), neutron star EOS constraints Legred et al. (2021), and semi-analytical ejecta fits Krüger and Foucart (2020), we found that BNS mergers following the sGRB-inferred delays cannot reproduce the observed disk-star [Eu/Fe] trend. An additional enrichment channel that tracks the SFR is therefore required, consistent with similar conclusions in the literature Côté (2019); Kobayashi et al. (2020, 2023). We note that some studies on the metal-poor halo stars and dwarf galaxies have emphasized the importance of delayed r-process sources in ultra-low-metallicity environments (e.g., Duggan et al. (2018); Tarumi et al. (2021)). These works probe the earliest phases of chemical enrichment, whereas our analysis focuses on the [Eu/Fe] evolution of Milky Way disk stars at higher metallicities.
Given the strong evidence for a prompt enrichment channel, we investigate whether two commonly proposed merger-based candidates—fast-merging BNS systems and NSBH mergers—can plausibly supply the needed early -process production. Fast-merging BNS may originate from dynamical interactions in dense stellar environments such as globular clusters or galactic nuclei Samsing et al. (2018); Rodriguez et al. (2018a), and are often hypothesized to merge promptly or eject more -process material Vigna-Gómez et al. (2018). NSBH mergers have similarly been proposed as efficient ejectors under certain mass and spin configurations Foucart et al. (2018); Mapelli et al. (2018). Both channels have been suggested as possible solutions to the early-enrichment problem, yet it remains unclear whether either can in fact produce the rates and yields required by the observed [Eu/Fe]–[Fe/H] trends. This is the primary question we resolve here.
Using a likelihood-based inference framework proposed in Chen et al. (2025), we assess whether NSBH and fast-merging BNS channels can contribute sufficiently to match the observed evolution of [Eu/Fe] versus [Fe/H], and whether the required conditions are astrophysically plausible given current observational constraints. While the point-estimate studies have shown that standard BNS mergers are unlikely to dominate early Galactic r-process enrichment, such estimates rely on representative values for merger rates, delay times, and ejecta yields. In reality, the r-process yield of compact-binary mergers depends sensitively on the component masses, spins, and equation of state, and the underlying merger population is only weakly constrained by current observations. A likelihood-based inference framework is therefore required to propagate these uncertainties consistently and to determine whether any plausible region of this multi-dimensional parameter space allows merger-based channels to dominate early enrichment. Moreover, point estimates cannot quantify how extreme the required conditions are—for example, whether an additional prompt channel must contribute a modest fraction or more than 95% of the total r-process budget. By marginalizing over population, yield, and observational uncertainties, our approach provides a quantitative assessment of whether fast-merging BNS or NSBH systems can plausibly serve as the missing prompt enrichment source.
The remainder of this paper is organized as follows. In Methods, we outline the chemical evolution model and describe our inference framework. In Results, we discuss results for the three scenarios: fast-merging BNS assumed to be GW-detectable, fast-merging BNS assumed to be GW-undetectable, and NSBH as a secondary channel. Finally, we summarize our findings and discuss their implications for the origin of -process elements in the Milky Way.
II Methods
II.1 Chemical Evolution Framework with Multiple -process Channels
Our chemical evolution model, incorporating multiple candidate -process enrichment channels, builds directly upon the one-zone prescription employed in Chen et al. (2025), which in turn follows the approach of Siegel et al. (2019b). In this framework, the Milky Way is treated as a single, well-mixed reservoir of interstellar medium (ISM). The model evolves the abundances of heavy elements over cosmic time by integrating the -process yields from astrophysical events convolved with their respective rate over time. If the astrophysical events were delayed relative to the SFR, we adopt a delay-time model, which is a power law distribution with minimum delay time () and slope () as parameters (e.g. Eq. 2 in Zevin et al. (2022)). For a given SFR, delay-time model, and yields, the model tracks the enrichment history in both and iron-peak elements, as well as -process elements such as Eu.
Key enrichment channels in the baseline model include core-collapse supernovae (CCSNe) and Type Ia supernovae (SNe Ia), mainly contributing to and iron enrichment. The -process enrichment is primarily attributed to delayed BNS mergers, but the model is flexible enough to include additional candidate channels such as fast-merging BNS systems and NSBH mergers, as we explore in this work. Each channel is parameterized by a local event rate, per-event -process yield or simply yield, and an evolutionary prescription: it either tracks the SFR directly or is obtained by convolving the SFR with a channel-specific delay–time distribution. For comparison with the observed [Eu/Fe]–[Fe/H] trends, we convert the total -process yield of each event into a Eu yield by assuming a fixed -process abundance pattern and adopting the solar -process Eu mass fraction. Thus, unless explicitly stated otherwise, “yield” refers to the total -process mass, and Eu yields are obtained as a fixed fraction of this quantity. In all cases, the underlying SFR is taken to be the cosmic star-formation history of Madau and Fragos (2017).
The evolution of the ISM, in addition to all the contributions from various channels, is also governed by a mass-loss term due to star formation and Galactic outflows, which scales with the SFR and which was calibrated in previous work Hotokezaka et al. (2018). For CCSNe and SNe Ia, we follow the same rates and yields for iron production as in Siegel et al. (2019b). Thus, the only free parameters in our model are for the -process sources. Given an arbitrary choice of parameters describing each -process channel, the model produces predictions for the time evolution of abundance ratios such as [Fe/H] and [Eu/Fe]. It is this [Eu/Fe] versus [Fe/H] abundance evolution track that we compare with the observed stellar abundance distributions to constrain possible -process enrichment channels.
II.2 Likelihood Model and Bayesian Inference Framework
The stellar abundance data of the disk stars we use are from Ref. Battistini and Bensby (2016). To quantify the agreement between model predictions and stellar abundance data, we adopt the likelihood formalism introduced in Ref. Chen et al. (2025). Each observed disk star is represented as a 2D Gaussian in ([Fe/H], [Eu/Fe]) space, centered on the measured values with widths set by observational uncertainties Battistini and Bensby (2016). We use only disk stars, as the one-zone chemical evolution model assumes a well-mixed ISM—an assumption more appropriate for the Galactic disk than for the halo. Halo stars, which formed earlier in more chemically inhomogeneous environments Deason and Belokurov (2024); Mackereth et al. (2019); Khoperskov et al. (2023); de Brito Silva et al. (2024), are therefore excluded.
The chemical evolution model described in previous subsection generates a time-ordered track in the same abundance space, and the likelihood for each star is computed by integrating its Gaussian distribution over the model track. The total likelihood is the product over all stars, evaluated numerically. We use this likelihood for Bayesian inference on the parameters governing each -process channel, including the local event rate, yield, and, where applicable, the minimum delay time () and delay-time distribution power-law index ().
For the local merger rates and r-process yields of BNS and NSBH systems, we begin with population inference results of component masses and local merger rates as obtained from the analysis of GWTC-4.0 The LIGO Scientific Collaboration, the Virgo Collaboration, the KAGRA Collaboration (2025). These population samples are propagated through the framework of Chen et al. (2021) to obtain joint constraints for the local merger rates and the ejecta masses. In this procedure, neutron-star radii and compactnesses are derived by combining the component masses with independent EOS constraints from joint GW and pulsar observations Legred et al. (2021). The resulting compactnesses are then used as inputs to NR-calibrated semi-analytical fitting formulae Krüger and Foucart (2020) to compute the dynamical ejecta and remnant disk masses with uncertainties as specified in Eq. (1) of Ref. Chen et al. (2025), which are subsequently combined to estimate the total ejecta mass. The resulting joint posterior distributions for the local merger rate and ejecta mass are considered as informative priors in our inference framework.
III Results
III.1 Current GW searches can detect fast-merging BNSs
We first examine whether a population of fast-merging BNS can account for the missing -process enrichment. In some extreme scenarios, the fast-merging BNSs may be difficult to search with current GW detection pipelines. For example, BNSs with high eccentricities may be missed. These systems are expected to form in dense stellar environments such as globular clusters or nuclear star clusters Samsing et al. (2014); Rodriguez et al. (2018b); Fragione et al. (2019), where close interactions can lead to highly eccentric binaries. The resulting rapid inspirals may exhibit significantly fewer in-band GW cycles, producing waveforms that are poorly matched to standard quasi-circular templates Chaurasia et al. (2018). This reduces their detection probability and may lead to systematic biases in parameter estimation Huez et al. (2025).
In this subsection, we discuss the scenario that fast-merging systems are detectable with the same efficiency as delayed BNS mergers. We will discuss the scenario that the fast-merging population is missed from current observations in the next subsection. While these are two extreme scenarios, the reality is likely to be in between the results we find from these two scenarios.
If the fast-merging population is detectable, current GW-inferred BNS merger rates should encompass both delayed and fast-merging systems. The total -process contribution from BNS mergers is then a mixture of both components. We parameterize the fraction of total BNS mergers that are fast-merging as
where is the total BNS rate, is the fast-merging BNS rate, and is the delayed BNS rate. We assume delayed BNS mergers follow an sGRB-informed delay-time distribution Zevin et al. (2022), while fast-merging BNSs are assumed to be SFR-tracking (i.e., zero delay time). Because of the differing evolution of rates for the two channels, the above rates and are functions of redshift. We then infer under this dual-BNS population scenario.
We assume that all BNS populations share the same neutron star EOS and follow the same prescriptions to calculate the -process yields given the component masses Krüger and Foucart (2020); Chen et al. (2025).
Our inference shows that reproducing the observed [Eu/Fe] trend in this scenario requires an exceptionally high fast-merging fraction. The posterior for in the local Universe () peaks near unity, with more than of BNS mergers inferred to be fast-merging (90% confidence lower limit). This fraction goes further up at higher redshifts, as a SFR-tracking channel is expected to contribute more frequently in the early Universe compared to delayed sources. In Figure 1, we show the corresponding volumetric rate as a function of redshift. In particular, the local () volumetric rate of the fast-merging BNS is Gpc while the delayed BNS rate remains at a lower level of Gpc, where the error bars correspond to 90% credible levels (The reader may refer to the Appendix and Fig. 4 for the posteriors of all inferred parameters.). Since this analysis assumes that GW searches are sensitive to both channels, the sum of the two channels is limited by the GW-inferred local BNS rate (). It is not the case for our analysis in the next subsection.
These findings suggest that, even with zero delay time, the fast-merging BNSs have to show a near-total dominance in the entire BNS population in order to account for the -process abundance evolution in the Milky Way—specifically the negative slope in [Eu/Fe] versus [Fe/H] for the disk stars. This scenario poses a strong tension with the current understanding of fast-merging BNS populations (Beniamini and Piran, 2019; Belczynski et al., 2018; Giacobbo and Mapelli, 2018), and hence challenges the idea that they are the primary explanation for Galactic -process enrichment.
III.2 Current GW searches miss fast-merging BNSs
In this scenario, we consider the possibility that fast-merging BNSs constitute a distinct subpopulation that is not captured by current GW searches. We note that although eccentric systems are good examples, our study applies to any scenario in which fast-merging BNSs are missed.
To model this population, we assume that delayed BNS mergers follow the standard GW-inferred rate and sGRB-informed delay-time distribution Zevin et al. (2022), as in the previous subsection. We then introduce an additional -process channel composed of GW-undetectable fast-merging BNS systems. This additional population is assumed to track the SFR.
The contribution from this channel is parameterized using a multiplicative enhancement factor , such that the total rate from undetectable fast-merging BNS mergers is:
where is the rate of delayed BNS mergers. The local () value of is inferred from GW observations. Again, the rates and are functions of redshift.
Our analysis finds that, in order to explain the observed [Eu/Fe] versus [Fe/H] trend (negative slope for disk stars, as mentioned before) with only delayed and undetected fast-merging BNSs, the multiplicative factor is estimated to be in the local Universe (90% confidence lower limit), while it further rises up to at and at . Equivalently, the fast-merging BNS rate would be Gpc in the local Universe (Figure 2, 90% confidence lower limit), and goes up to Gpc at and Gpc at . This implies that the Milky Way would require roughly 17 times (median value) as many fast-merging BNSs as delayed ones at ; this outcome again challenges the current understanding of the delayed and fast-merging BNS rates (Beniamini and Piran, 2019; Belczynski et al., 2018; Giacobbo and Mapelli, 2018) (see the Appendix and Fig. 5 for the posteriors of all inferred parameters.).
From the results in the previous and this subsection, we conclude that fast-merging BNSs, even with SFR-tracking rates, are difficult to resolve the -process enrichment discrepancy without invoking an extremely high occurrence rate relative to the delayed systems.
The posterior distributions for and , and the corresponding fast-merging BNS rates derived above quantify the event rates that would be required for fast-merging systems to reproduce the observed Galactic [Eu/Fe] trends within our chemical evolution model. They do not, by themselves, represent the probability that such fast-merging populations exist in nature. A fully self-contained Bayesian hypothesis test of fast-merging–dominated enrichment would require well-calibrated external priors on the intrinsic fast-merging BNS rate, for example from population synthesis, short-GRB statistics, or detailed modeling of dense stellar environments. At present, we lack any widely-accepted quantitative constraints on the fast-merging subpopulation in a form that can be robustly incorporated into a second-stage Bayesian model comparison. Our assessment therefore rests on the tension between the exceptionally high fast-merging rates inferred here and current theoretical and observational expectations for compact-binary populations, rather than on a formal statistical exclusion based on externally imposed rate priors.
III.3 NSBH Mergers as a Secondary Channel
In this subsection, we assess whether the combined contributions of BNS and NSBH mergers can explain the observed Eu enrichment in the Milky Way. To that end, we construct a joint four-dimensional distribution over the local merger rates and yields of BNS and NSBH systems, following the methodology of Chen et al. (2021) and using the fitting formulae from Krüger and Foucart (2020) for the yield with uncertainties. We use this distribution as priors for our chemical evolution modeling.
As mentioned earlier, the BNS delay-time distribution parameters (, ) are sampled from sGRB-informed constraints Zevin et al. (2022). The delay time for NSBH mergers is uncertain. It could follow the sGRB delay time, or have a completely different delay-time distribution. If NSBH mergers promptly merge with no delay relative to the SFR, it would contribute most to rectify the difference between the delayed BNS-only evolution track and the [Eu/Fe] versus [Fe/H] observations, as motivated by Chen et al. (2025). We start with this assumption.
We find that even under optimistic delay-time distribution assumptions to maximize the contribution from NSBH, the combined BNS and NSBH channels fail to reproduce the observed [Eu/Fe] trends.
To allow for residual enrichment not captured by BNS and NSBH channels, we then add a third, phenomenological channel whose rate also tracks the SFR. Its contribution is parameterized as a fraction of the total -process enrichment, as its rate and yield are unknown. This component captures any dominant but currently unidentified source of early-time Eu production.
Figure 3 shows the inferred (rate yield) fraction integrated over cosmic history until present day, from the BNS (blue), NSBH (orange), and third (green) channels. The BNS contribution is consistent with the findings of Chen et al. (2025), while the NSBH contribution remains marginal. In contrast, the third, phenomenological SFR-tracking channel dominates the posterior. Quantitatively, at 90% confidence, more than 90% of the contribution must be from the unknown third channel while the combined contribution from BNS and NSBH is limited to below 10% (The reader may refer to the Appendix and Fig.. 6 for the posteriors of all inferred parameters.). These results underscore that BNS and NSBH mergers alone cannot account for the observed [Eu/Fe] versus [Fe/H] trend, and that a substantial additional source of -process elements is required.
IV Discussion and Conclusion
We have investigated whether fast-merging BNSs or NSBH mergers can resolve the -process enrichment discrepancy identified previously Côté (2019); Cowan et al. (2021); Kobayashi et al. (2023); van de Voort et al. (2015); Siegel et al. (2019b); Chen et al. (2025), where BNS mergers fail to reproduce the observed Eu trends—the negative slope in [Eu/Fe] versus [Fe/H]—in the Milky Way disk. Across all three scenarios explored—GW-detectable fast-merging BNSs, GW-undetectable fast-merging BNSs, and NSBH mergers—we find that, under current theoretical and observational expectations for compact-binary populations, these channels are unlikely to serve as the dominant source of -process enrichment. This conclusion reflects the tension between the event rates required by our chemical-evolution inference and existing expectations for fast-merging and NSBH populations.
Although the sGRB host observations used to infer the delay-time distribution Zevin et al. (2022) may suffer from selection biases that favor nearby events, potentially missing fast-merging systems occurring at higher redshifts, our analysis shows that the fast-merging population must be significantly larger than the delayed population—even in the local Universe. This result raises several major challenges for both theories and observations:
1.
What is the coherent theoretical framework that explains why fast-merging BNSs form significantly more efficiently than delayed systems across the Milky Way history?
2.
Even if high-redshift hosts of sGRBs are missed, why do nearby sGRBs predominantly exhibit long delay times, in apparent contradiction with the presence of a large fast-merging population?
3.
If instead fast-merging BNSs do not produce observable sGRBs, but do synthesize substantial quantities of -process elements (possibly accompanied by significant amount of kilonovae), what distinguishes their counterpart production mechanisms from those of delayed BNS mergers?
4.
If a large number of fast-merging BNSs exist that are not detectable in GWs, why have current GW search pipelines—both modeled and unmodeled Abac et al. (2025)—not identified nearby events with high signal-to-noise ratios, especially given that some of these should be detectable even if existing waveform models are suboptimal?
We emphasize that, although the fast-merging BNS and NSBH populations may not precisely follow the assumed star formation rate, may have different yields from those estimated with semi-analytical fit, may have mass distributions that differ from those inferred from GW observations, and/or may not adhere to the assumed -process abundance pattern—all of which could affect their event rates or -process yields—our analysis still indicates that an exceptionally large contribution from these channels is required across cosmic history (see the required fractions and volumetric rates in the Results section). This poses a substantial challenge to constructing a consistent theoretical framework. For instance, to reduce the event rates of fast-merging BNSs or NSBHs to a level comparable to that of delayed BNSs, their yield would need to be increased by more than an order of magnitude. This requires a scenario that boosts their yields by an order of magnitude without increasing the delayed BNS yield. Similarly, even if the early-Universe evolution of these populations differs from our assumptions, they would still need to occur much more frequently than delayed BNSs to reproduce the observed trends in stellar metallicity. To explore a very different SFR, we repeated our analysis assuming a constant SFR and found that our conclusions remain unchanged (see Appendix-B and Fig. 7 for more details).
While we have assumed identical r-process yields for delayed and fast-merging BNSs, this assumption is not required. In a more general scenario, the yields from these two channels may differ. In particular, systems classified as fast-merging may preferentially form on highly eccentric orbits, perhaps leading to more violent encounters than quasi-circular binaries and potentially producing a larger mass (25-30% more) of neutron-rich ejecta, as suggested by previous studies Papenfort et al. (2018); Foucart et al. (2024). To explore this possibility, we repeat our fast-merging BNS analyses assuming a yield that is twice that of delayed BNSs (100% more). Even under this optimistic assumption, our results indicate that fast-merging BNSs must still dominate the total merger rate: they must contribute more than 94% of the total rate in the scenario where GW searches are sensitive to fast mergers, and require a local-universe rate enhancement of when GW searches are not sensitive to fast mergers.
We note that this conclusion is also supported by independent astrophysical considerations regarding the environments in which fast-merging BNS systems are expected to form. Such systems are commonly associated with dense stellar environments, including globular clusters and nuclear star clusters. In the Milky Way, globular clusters are predominantly old, metal-poor, and reside in the halo, with star formation ceasing well before the formation of the thin disk considered here, making efficient enrichment of disk stars unlikely. Nuclear star clusters, while more metal-rich, are spatially concentrated near the Galactic center and are similarly disfavored as major contributors to the r-process content of the solar neighborhood. These considerations provide complementary physical support for our conclusion that fast-merging BNSs are unlikely to dominate r-process enrichment in the Galactic thin disk.
Our findings not only reinforce the conclusions of Chen et al. (2025), but also pose substantial challenges to two leading candidates—fast-merging BNS and NSBH systems—as dominant -process enrichment channels. These results help narrow down the viable -process sources in the Milky Way and strongly suggest that non-merger pathways are needed to explain the observed abundances.
Acknowledgements.
The authors would like to thank Floor Broekgaarden, Wen-fai Fong, Francois Foucart, Erika Holmbeck, and Alexander Ji for very helpful discussions. M.S. is supported by the Weinberg Institute for Theoretical Physics at the University of Texas at Austin. H.-Y.C. is supported by the National Science Foundation under Grant PHY-2308752 and Department of Energy Grant DE-SC0025296. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.Appendix A Additional details for all three cases
In this appendix, we present additional plots that shed light on further details of each of the three analyses discussed in the main text. Figure 4 shows corner plots of the posterior distributions for all model parameters: the total BNS rate (including both delayed and fast–merging channels); the ejecta mass (i.e., the per-event -process yield, assumed common to both channels); the delay–time parameters and of the delayed BNS channel; and the fraction , which specifies the portion of assigned to the fast–merging channel required to reproduce the observed abundance trends of Milky Way disk stars Battistini and Bensby (2016). The large panel in the upper right shows the 90% credible regions (orange) in the – plane implied by the posterior samples; observed abundances from Battistini and Bensby (2016) are overplotted as blue points.
Fig. 5 presents the similar details for the analysis discussed in Fig. 2, in which refers exclusively to the delayed BNS channel. The ejecta mass is again taken to be common, and the fast component is encoded via a multiplicative factor such that .
Finally, Fig. 6 summarizes the analysis in which we include both BNS and NSBH contributions with an agnostic third channel to capture any residual enrichment not accounted for by them. The sampled parameters are the rate times mass ejecta products and , the BNS channel delay-time parameters , and a fractional contribution parameter that encodes the third channel’s share of the total mass–injection rate (we take with ). The large panel at upper right again shows the posterior–predicted credible regions in the – plane with the same observational sample overplotted.
Appendix B Impact of star formation history models
In this section, we assess the sensitivity of our results to the assumed SFR model. We repeat the full inference under a constant SFR—an extreme limiting case—and compare it to the cosmic SFR used in the main text. The three panels in Fig. 7 summarize the outcome for each secondary–channel scenario considered in the paper: (i) fast–merging BNS assumed detectable by current GW searches; (ii) fast–merging BNS assumed missed by current searches; and (iii) NSBH mergers. Across all three cases the posterior shifts between the two SFR assumptions are modest, and the qualitative conclusions of the paper are unchanged: neither fast–merging BNS nor NSBH mergers can account for the Milky Way’s -process enrichment. Bayesian model selection decisively prefers the cosmic SFR in every scenario, with logarithmic Bayes factors , , and , respectively, for the three cases. This justifies adopting the cosmic SFR in the main analysis and demonstrates that our main findings are robust to the SFR choice.